OpenStudy (anonymous):

Solve using Quadratic Formula 2x^2 + x - 6 = 0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

i solved it and got the wrong answer

OpenStudy (anonymous):

could you show the steps?

hartnn (hartnn):

what did u take the values of a,b,c as ?

hartnn (hartnn):

are these the values u took? a=2,b=1,c=-6 ?

hartnn (hartnn):

STEPS: \(\huge{x=\frac{-1 \pm \sqrt{1-4(2)(-6)}}{2(2)}}=\frac{-1 \pm \sqrt{(1+48)}}{2(2)}=\\ \huge\frac{-1 \pm7}{2(2)}=\frac{6}{4}\quad or \quad\frac{-8}{4}=1.5\quad or \quad -2 \)

OpenStudy (anonymous):

ohhhkay i get it. I wanted to stop at the 3rd step, but you have to actually add or subtract it. Thank you!

hartnn (hartnn):

welcome :)